A spring scale hung from the ceiling stretches by 5.3 cm when a 1.8 kg mass is hung from it. The 1.8 kg mass is removed and replaced with a 2.5 kg mass.
Answer #1

Apply Hooks law

F = kx

k = F/x

= mg/x

= 1.8 (9.8)/5.3* 10^-2)

=332.83 N/m

when m = 2.5 kg

x= F/k = 2.5(9.8)/332.83

= 7.36 cm